The sum of two numbers is twenty. Three times the smaller is equal to twice the larger. Find the two numbers.
Answer by tutorcecilia(2144)

Translate from English to Math:
x+y=20 [The sum of two numbers is twenty]
3x=2y [Three times the smaller is equal to twice the larger]
.
x=20-y [Solve for x]
3(20-y)=2y [Substitute (x=20-y) into the second equation]
60-3y=2y [Solve for y]
60=5y
60/5=y
12=y
.
x+y=20 [Plug (y=12) back into either of the original equations]
x+12=20 [solve for x]
x=8
.
Check by plugging the values of x and y into either equation.